# Rewarding Graph Reasoning Process makes LLMs more Generalized Reasoners

### v1

@@ Abstract.
Despite significant advancements in Large Language Models (LLMs), developing advanced reasoning capabilities in LLMs remains a key challenge. Process Reward Models (PRMs) have demonstrated exceptional promise in enhancing reasoning by providing step-wise feedback, particularly in the context of mathematical reasoning. However, their application to broader reasoning domains remains understudied, largely due to the high costs associated with manually creating step-level supervision. In this work, we explore the potential of PRMs in graph reasoning problems - a domain that demands sophisticated multi-step reasoning and offers opportunities for automated step-level data generation using established graph algorithms. We introduce GraphSilo, the largest dataset for graph reasoning problems with fine-grained step-wise labels, built using automated Task-oriented Trajectories and Monte Carlo Tree Search (MCTS) to generate detailed reasoning steps with step-wise labels. Building upon this dataset, we train GraphPRM, the first PRM designed for graph reasoning problems, and evaluate its effectiveness in two key settings: inference-time scaling and reinforcement learning via Direct Preference Optimization (DPO). Experimental results show that GraphPRM significantly improves LLM performance across 13 graph reasoning tasks, delivering a 9% gain for Qwen2.5-7B and demonstrating transferability to new graph reasoning datasets and new reasoning domains like mathematical problem-solving. Notably, GraphPRM enhances LLM performance on GSM8K and MATH500, underscoring the cross-domain applicability of graph-based reasoning rewards. Our findings highlight the potential of PRMs in advancing reasoning across diverse domains, paving the way for more versatile and effective LLMs. Models and data are available at https://huggingface.co/GraphPRM.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
@@ 1. Introduction
2. Preliminaries and Related Works
2.1. LLMs for Graph Reasoning
2.2. Reward Models for Reasoning Problem
Reasoning - the capacity to derive new insights, tackle challenges, and reach conclusions from given information - stands as a core capability of LLMs. Researchers have proposed various methods to improve or elicit the reasoning ability of LLMs (e.g., mathematical reasoning), which can be broadly divided into three groups: pre-training (OpenAI, 2023; AI@Meta, 2024; Team, 2024b, a), fine-tuning (Yu et al., 2024; Luo et al., 2023; Yue et al., 2024; Chen et al., 2024c) and prompting (Imani et al., 2023; Wang et al., 2023c; Chen et al., 2024b). Except for directly improving and eliciting the reasoning potential of LLMs, researchers have been dedicated to boosting LLM reasoning results via an extra verifier for selecting the best answer from multiple decoded candidates. There are two primary types of verifiers: the Outcome Reward Model (ORM) (Cobbe et al., 2021b) and the Process Reward Model (PRM) (Lightman et al., 2024; Wang et al., 2024b; Zheng et al., 2024).
@@ 3.1.4. GraphSilo Statistics
This extensive dataset provides a strong foundation for training step-wise PRM and evaluating LLMs on various GCP tasks. As shown in Table 2, our dataset is the largest one to date containing CoT solutions, with 118,189 samples and 394,165 step-wise labels. It is also the only dataset that includes step-level supervision, offering a valuable resource for graph reasoning research. The detailed comparison with previous related benchmarks is shown in Table 2.
@@ Appendix B Additional Experimental Results
In this section, to further analysis our proposed GraphPRM, we evaluate GraphPRM by addressing the following question: ”How does GraphPRM perform on other GCP-specific backbones?”
Performance of GraphPRM on GraphWiz model In addition to our proposed GraphSilo, we conduct experiments on GraphWiz dataset with Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct and LLaMA-3.1-8B-Instruct in Section 4.2. Besides above, we further evaluate the transferbility of GraphPRM on other GCP-specific model like GraphWiz. As shown in Figure 10, GraphPRM consistently improve LLMs’ performance on GraphWiz dataset with more than 5% improvement.

### v2

@@ Abstract.
Despite significant advancements in Large Language Models (LLMs), developing advanced reasoning capabilities in LLMs remains a key challenge. Process Reward Models (PRMs) have demonstrated exceptional promise in enhancing reasoning by providing step-wise feedback, particularly in the context of mathematical reasoning. However, their application to broader reasoning domains remains understudied, largely due to the high costs associated with manually creating step-level supervision. In this work, we explore the potential of PRMs in graph reasoning problems - a domain that demands sophisticated multi-step reasoning and offers opportunities for automated step-level data generation using established graph algorithms. We introduce GraphSilo, the largest dataset for graph reasoning problems with fine-grained step-wise labels, built using automated Task-oriented Trajectories and Monte Carlo Tree Search (MCTS) to generate detailed reasoning steps with step-wise labels. Building upon this dataset, we train GraphPRM, the first PRM designed for graph reasoning problems, and evaluate its effectiveness in two key settings: inference-time scaling and reinforcement learning via Direct Preference Optimization (DPO). Experimental results show that GraphPRM significantly improves LLM performance across 13 graph reasoning tasks, delivering a 9% gain for Qwen2.5-7B and demonstrating transferability to new graph reasoning datasets and new reasoning domains like mathematical problem-solving. Notably, GraphPRM enhances LLM performance on GSM8K and MATH500, underscoring the cross-domain applicability of graph-based reasoning rewards. Our findings highlight the potential of PRMs in advancing reasoning across diverse domains, paving the way for more versatile and effective LLMs. Codes and data are available at https://github.com/GKNL/GraphPRM.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737109isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15553493.
@@ 1. Introduction
2. Preliminaries and Related Works
2.1. LLMs for Graph Reasoning
2.2. Reward Models for Reasoning Problem
Reasoning - the capacity to derive new insights, tackle challenges, and reach conclusions from given information - stands as a core capability of LLMs. Researchers have proposed various methods to improve or elicit the reasoning ability of LLMs (e.g., mathematical reasoning), which can be broadly divided into three groups: pre-training (OpenAI, 2023; AI@Meta, 2024; Team, 2024b, a), fine-tuning (Yu et al., 2024; Luo et al., 2023; Yue et al., 2024; Chen et al., 2024c; Wu et al., 2025) and prompting (Imani et al., 2023; Wang et al., 2023c; Chen et al., 2024b). Except for directly improving and eliciting the reasoning potential of LLMs, researchers have been dedicated to boosting LLM reasoning results via an extra verifier for selecting the best answer from multiple decoded candidates. There are two primary types of verifiers: the Outcome Reward Model (ORM) (Cobbe et al., 2021b) and the Process Reward Model (PRM) (Lightman et al., 2024; Wang et al., 2024b; Zheng et al., 2024).
@@ 3.1.4. GraphSilo Statistics
This extensive dataset provides a strong foundation for training step-wise PRM and evaluating LLMs on various GCP tasks. As shown in Table 2, our dataset is the largest one to date containing CoT solutions, with 118,189 samples and 394,165 step-wise labels. It is also the only dataset that includes step-level supervision, offering a valuable resource for graph reasoning research. We report detailed dataset statistics in Table 9.
@@ 4. Experiments
More results and analysis about GraphPRM on different models and graph densities are presented at Appendix B.
@@ 4.2. Performance of GraphPRM (RQ1)
Comparison against ORMs/Math-PRMs As GraphPRM is the first PRM specific for GCP tasks, to further demonstrate the effectiveness of GraphPRM compared to ORM, we train an additional G-ORM on GraphSilo without step-wise labels for GCP tasks (only use outcome reward for supervision). As reported in Table 4, we can observe that though G-ORM outperforms SC, it lags behind GraphPRM on both in-domain and out-domain graph reasoning tasks, which further prove the effectiveness of GraphPRM. Beyond this, we also investigate the performance of representative math-specific PRMs on GraphSilo, and Table 4 shows that math-specific PRMs cannot easily transfer to graph reasoning tasks. We conclude this phenomenon for the following reasons: 1) Graph reasoning tasks involve diverse reasoning patterns (e.g., topological, logical, and computational reasoning) that inherently require an understanding of graph structures, which is less pronounced in mathematical problem-solving. 2) The difficulty of graph problems often scales exponentially with the size and complexity of the graph.
Methods In-domain Out-domain Average
Self-Consistency 37.03 56.67 41.56
G-ORM 37.40 56.33 41.77
G-ORM + Self-Consistency 37.77 57.00 42.21
Math-Shepherd-7B (Wang et al., 2024b) 36.23 54.89 40.54
Math-PSA-7B (Wang et al., 2024a) 37.30 56.44 41.72
GraphPRM (Best-of-N) 38.30 58.44 42.95
GraphPRM (Beam Search) 41.57 57.67 45.28
Table 4. Comparison among GraphPRM, G-ORM and math-specific PRMs on GraphSilo. The policy model is Qwen2.5-7B-Instruct. The generation budget during verification is set to 8. The best results are in Purple.
@@ 4.3. Transferability on Math Problems (RQ2)
Comparison against Math-PRMs on math datasets We further compare GraphPRM with two representative math-specific PRMs: Math-Shepherd-Mistral-7B (Wang et al., 2024b) and MATH-PSA-7B (Wang et al., 2024a). Results reported in Table 6 shown that GraphPRM yield strong and competitive performance on math reasoning benchmarks. With the observation in Table 4 that math-specific PRMs cannot easily transfer to graph reasoning tasks, it further demonstrates the uniqueness and effective transferability of GraphPRM.
Methods GSM8K MATH
Best-of-N Beam Search Best-of-N Beam Search
Self-Consistency 89.08 89.08 75.40 75.40
Math-Shepherd-7B 90.09 89.08 75.90 76.00
Math-PSA-7B 90.50 90.25 76.30 75.40
GraphPRM 89.92 90.54 76.00 76.30
Table 6. Comparison among GraphPRM and math-specific PRMs on math datasets. The policy model is Qwen2.5-Math-1.5B-Instruct. The generation budget during verification is set to 8. The best results are in Purple.
@@ 5. Conclusion
6. Acknowledgements
We would like to thank all the anonymous reviewers and area chairs for their insightful and valuable comments. We also thank the support of the 12th Baidu Scholarship. This research is supported by National Key Research and Development Program of China Grant No.2023YFF0725100 and Guangdong S&T Program C019.
@@ Appendix B Additional Experimental Results
In this section, to further analysis our proposed GraphPRM, we evaluate GraphPRM by addressing the following extra research questions:
  • RQ7: How does GraphPRM perform on other GCP-specific backbones?
  • RQ8: How does graph density affect the performance of GraphPRM on LLMs during inference time?
B.1. Performance of GraphPRM on GraphWiz model (RQ7)
In addition to our proposed GraphSilo, we conduct experiments on GraphWiz dataset with Qwen2.5-7B-Instruct, Qwen2.5-Math-7B-Instruct and LLaMA-3.1-8B-Instruct in Section 4.2. Besides above, we further evaluate the transferbility of GraphPRM on other GCP-specific model like GraphWiz. As shown in Figure 10, GraphPRM consistently improve LLMs’ performance on GraphWiz dataset with more than 5% improvement.
@@ B.1. Performance of GraphPRM on GraphWiz model (RQ7)
B.2. Impact of Graph Density (RQ8)
Graph density can be a crucial factor influencing the performance of GraphPRM on graph reasoning tasks. We regard Edge Density $\left(x=\frac{2|E|}{n(n-1)}\right)$ as the reflection of graph density, and we construct new test set with three-level density: Low ($0<x\leq 0.33$), Middle ($0.33<x\leq 0.67$) and High ($0.67<x\leq 1$). Results with different graph densities are shown in figure 11. It can be observed that accuracy decreases as graph density increases, but the model with GraphPRM consistently outperforms that with Self-Consistency, demonstrating its robustness across graphs with different densities. Besides, as graph density increases, the performance gained via GraphPRM (Best-of-8) decreases. We conclude this is the reason that: Graphs with higher density contain more invalid edges or noisy edges, bringing more challenges for GraphPRM to identify essential steps. This further highlights the importance of employing GraphPRM to supervise step-wise graph reasoning.
Figure 11. Best-of-N Performance of GraphPRM on GraphSilo with different graph densities. The policy model is Qwen2.5-7B-Instruct and the verification is based on 8 candidates.